OpenStudy (anonymous):

State whether the sequence is arithmetic. If so, state the common difference. 1.8, 3.6, 7.2, 14.4, ... Answer 1.8 -1.8 not arithmetic 16.2

OpenStudy (anonymous):

Here, find t3-t2 and t2-t1

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

If t3-t2=t2-t1 then arithmetic with common difference=t3-t2

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So C ?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

it is C